OpenStudy (anonymous):

If the coordinates at the endpoint of P(t) are (-1/3, 2 sq rt 2/3). what would be the coordinate at the endpoint of each of the following. a) P(t+pi)

OpenStudy (anonymous):

@aaronq

OpenStudy (adi3):

wht grade is that

OpenStudy (anonymous):

10

OpenStudy (adi3):

@math92130

OpenStudy (adi3):

ohh ok got it

OpenStudy (anonymous):

how do i solve it

OpenStudy (adi3):

give me a minute @math

OpenStudy (adi3):

@math92130

OpenStudy (michele_laino):

is P(t) as below? \[P\left( t \right) = \left( { - \frac{1}{3},\;2\sqrt {\frac{2}{3}} } \right)\]